# CHIMERA: A Knowledge Base of Idea Recombination in Scientific Literature

### v3

@@ 4.1 Recombination extraction evaluation
Generally, fine-tuning Mistral-7B using our data obtains the best performance in all subtasks. We observe that entity and relation extraction appear more challenging than classification for both humans and SOTA LLMs. However, humans still significantly outperform automatic extraction approaches. We present error analysis in Appendix F. Our findings indicate that focusing on a smaller portion of the recombination extraction task is not necessarily easier than performing it end-to-end, as seen in the lower performance of abstract classifiers, and discuss this point further in Appendix C.
4.2 Knowledge base analysis
@@
Figures 3(a) and 3(a) present the predominant domain pairs for inspiration and blend relations in CHIMERA (with frequency above the 0.9 quantile). Inspirations display a larger selection of domains than blends. We also observe that blends connect the same or similar domains, while inspirations are often between different and further domains. Of note is the volume of inspiration drawn from brain-related sources, such as cognitive science and q-bio.nc. A possible explanation might be that many of our arXiv categories of interest are related to machine learning, where the human brain historically serves as a general source of inspiration. Table 14 in Appendix H presents the same information in tabular format for better readability.
@@
We next analyze how different fields draw inspiration from each other. Figure 3(c) shows the top 10% cross-domain inspiration sources for three prevalent domains in the graph: cs.RO (Robotics), cs.CV (Computer Vision) and cs.CL (Computation and Language). We observe that while some sources of inspiration (like cognitive-science) are commonly shared across related fields, domains may draw inspiration from unique sources (e.g., from zoology to cs.RO). Examples of these interdisciplinary inspirations can be found in Appendix G, Table 13.
Figure 4 shows the percentage of target nodes in domains drawing inspiration from cs.CL (Computation and Language) over the past five years. We observe two trends: a decrease in intra-domain inspiration (where cs.CL concepts inspire other cs.CL concepts), and an increase in cs.CV (Computer Vision) concepts drawing inspiration from cs.CL.
5 Recombination Prediction
This section gives an example of a possible use case of the CHIMERA knowledge base. Using this data, we train supervised models that learn how to recombine concepts for predicting new scientific ideas. This process is illustrated in Figure 2.
5.1 Problem Definition
@@
Table 5: Prediction data splits. We divide our data by the publication years associated with each query (training and validation sets $<$ 2024, test set $\geq$ 2024) to avoid contamination.
Figure 2 shows an example of the inputs and outputs of this task: given a context string ("Recent advancements in video generation have struggled to model complex narratives…") and a query regarding the recombination of a graph entity ("What would be a good source of inspiration for video generation?") we aim to answer this question with an additional graph entity ("The concept of storyboarding…"). More formally, given a context string (e.g., a problem, experimental settings, goals), an entity $e$ and a recombination type $\tau$, we aim to find a different graph entity $e^{\prime}$ such that $(e,\tau,e^{\prime})$ is an edge in CHIMERA.
5.2 Recombination prediction
@@
Table 6: Recombination prediction results. MedR stands for "Median Rank". Using CHIMERA data to fine-tune the models improves the median rank by a factor of 10. Reranking the top-20 answers using RankGPT boosts the H@10 but slightly reduces other metrics (H@3,5 and MRR).
@@
We start by converting edges to pairs of queries and answers. The queries describe the task inputs: a single graph node, the edge recombination type, and a context string, which we extract from the corresponding abstract using GPT-4o-mini. Note that this process might leak information regarding the answer (the other graph node) into the query. Therefore, we follow it by applying GPT-4o-mini again to the result to identify leakages in the generated queries (see examples and additional implementation details for this step in Appendix I). We discard bad query-answer pairs (approximately $22\%$ of the pairs, mostly due to leaks) and divide the remaining pairs into data splits based on the publication year. Our test set includes all pairs associated with papers published after 2024. Table 5 shows a summary of the resulting data splits.
@@
We experiment with zero-shot and finetuned retrievers based on encoders trained before the test set cutoff year (2024). We next explore applying a GPT-4o-based reranker Sun et al. (2023) to the top 20 retrieved results to improve our predictions further. The GPT-4o data cutoff is October 2023, meaning the reranker is also unfamiliar with our test set. Appendix J provides additional implementation details for the prediction baselines.
@@ 5.3 Prediction results
We present our results in Table 6. We observe that fine-tuning greatly helps to improve retrievers, decreasing the median rank of the gold answer by one order of magnitude. The last row of Table 6 shows results for applying RankGPT Sun et al. (2023) with GPT-4o over the top-20 results of the best-performing retriever (all-mpnet-base-v2finetuned). While the reranker improves H@10, it degrades H@k for $k=3,5$ and MRR values. We hypothesize the reranker might lower the gold answer rank if there are many possible answers.
@@
We present a human evaluation study exploring our ideation approach compared to other baselines. We recruit three volunteers with verifiable research experience (the authors of at least one research paper) and assign them examples based on their area of expertise. The examples are inspiration queries from our test set, along with inspiration source suggestions from different baselines: (1) Ours: our method, including reranking (2) Gold: the gold answer, (3) Random: a random test-set node, (4) GPT-4o: a suggestion generated by GPT-4o, (5) ZS-CHIMERA: a zero-shot prediction model using our test nodes as candidates, and (6) ZS-SciERC: a zero-shot prediction model using candidates extracted from test set abstracts with the SciERC Luan et al. (2018) schema. Note that we use the highest ranked answer (k=1) for baselines returning a ranked list of candidates.
Figure 5: Researchers find our recombination suggestions almost as helpful as the gold answer in inspiring ideas, providing additional verification of our automated evaluation.
We request the annotators to rank baseline suggestions based on their helpfulness in inspiring interesting ideas. Figure 5 presents each baseline’s median and average rank over 70 annotated examples. Since the most helpful suggestions are ranked first, a lower median and average rank signifies a more helpful baseline. Our approach receives a similar median and average rank as the gold answer, and annotators prefer it to all other baselines. This gives as an additional, complementary signal to the automatic evaluation, showing that our recombination prediction approach learns to create helpful recombinations. Appendix K presents additional details regarding the task and interface we used to conduct the study.
@@ Conclusions
We automatically mine the scientific literature to create CHIMERA, a novel knowledge base spanning over 28K examples of how scientists blend concepts and draw inspiration from different areas. This knowledge base offers broad applications, and we demonstrate how it can be used to empirically study idea recombination across domains and to fine-tune models that predict new recombination directions that researchers find inspiring.
@@
While our information extraction model improves the quality of mined recombinations, it is still far from being perfect. Our qualitative error analysis shows the extraction model struggles to identify and extract more subtle recombination descriptions, and it still falls significantly short of human performance on the same task. Improving the extraction model remains a challenging and interesting direction for future work.
@@
One particularly challenging aspect of the recombination prediction task is the lack of a single correct response. Given a problem description, there are numerous ways to blend ideas and take inspiration that can lead to a novel, recombinant solution. This could lead to many false negatives, resulting in an overly pessimistic estimate of the models. We partially handle this through an additional human evaluation. However, given the high level of expertise required from the participants, the scope and thoroughness of such an evaluation are limited.
@@
We use models from the GPT-4o series for evaluation (judging entity span similarity), analysis (identifying entity’s scientific domain), and to enrich our data (generating a context string for the extracted recombinations). Since this work prioritises the extraction and prediction of recombinations, we only experiment with those models. We leave experimenting with a larger range of models for these tasks for future work.
Ethics Statement
To collect human-annotated recombination examples, we recruited crowdworkers via Upwork. Annotators were informed of the nature and purpose of the task and were compensated at an hourly rate of \$26–\$30. Additionally, three volunteers participated in our human study. No personal information about the annotators or volunteers is disclosed.
@@
To promote transparency and reproducibility, we release our code and model checkpoints. The collected data is shared under an open license to facilitate further research. We used AI assistants for grammatical corrections and code writing (e.g., GitHub Copilot).
Appendix A Recombination keywords
@@
Appendix B Extraction baselines implementation
Annotators’ agreement
We calculate the annotators’ agreement by treating one annotator’s work as the ground truth and the other as predictions. In addition to entity-level and relation-level agreement, we also measure agreement on the recombination presence - whether the text presents a recombination, regardless of type. We apply the soft entity and relation matching described in 4.1 to measure the relation and entity agreement. The agreement is computed over the $52$ documents shared by both annotators (approximately $10\%$ of all annotated data). We regard the annotators’ agreement as an approximation of human performance measurement over this task.
@@ Appendix B Extraction baselines implementation
We use Mistral-7B as the backbone for our recombination extraction baseline. We fine-tune the model using mistral-finetune [^5] https://github.com/mistralai/mistral-finetune on a single NVIDIA RTX A6000 48GB GPU over $500$ steps. The training was conducted using the default learning rate of $56.e-$ 5 and weight decay of $0.1$. We use a batch size of $1$ and a maximum sequence length of $4096$ tokens. mistral-finetune implements Low-Rank Adaptation of LLM (LoRA), a parameter efficient fine-tuning method Hu et al. (2021), which we use with the default rank of $64$. The evaluation uses the corresponding repository, mistral-inference [^6] https://github.com/mistralai/mistral-inference. We rerun the same experiment using Llama-3.1-8B as a backbone, using an additional 500 warm-up steps, a learning rate of $52e-$ 5 and a weight decay of $0.01$. Figure 6 presents the prompt for these experiments.
@@
In addition to fine-tuning LLMs on our data, we experiment with GoLLIE Sainz et al. (2023), a general IE model fine-tuned to follow any annotation guidelines in a zero-shot fashion. We apply GollIE-13B on our data, using a single NVIDIA RTX A6000 48GB GPU, 1-beam search, and limit the new token number to 128. GoLLIE is finetuned from CODE-LLaMA2, and receives guidelines in the form of data classes describing what objects and properties the model should extract. Figure 7 depicts the guidelines we used to test GoLLIE as an E2E recombination extraction model. In the rare cases where the model returns more than a single recombination type ($<10$), we select the first.
@@
We also experiment with GPT-4o in few-shot settings. We select 45 examples for each example type (blend, inspiration, not-present) from the training data (a total of 135). As Table 2 describes, the training set only has 45 inspiration examples (as opposed to $>100$ blend and not-present examples). $45$ is, therefore, the maximal number of examples per class we can sample while keeping the ICL set balanced. We run each experiment $5$ times, sampling a new set of few-shot examples in each, and report the average. Figure 8 presents the prompt for this experiment.
@@
To evaluate entity extraction, we compare our model against GPT-4o in few-shot settings and include $45$ cases per example type, similarly to the E2E experiment. To account for variability due to example selection, we run each experiment $5$ times, sampling a new set of few-shot examples in each, and report the average. The total cost of this process sums up to 50\$. The prompt template for this experiment is available on Figure 9.
@@
We experiment with non-generative approaches as well, and compare our model to a SciBERT Zhong and Chen (2021) based token classifier. The encoder uses a standard Hugging-Face implementation of SciBERT, which we train on a single NVIDIA RTX A6000 48GB GPU over 500 steps. We use a weight decay of $0.1$, a learning rate of $56.e-$ 5 and a batch size of $1$. We also experiment with PURE Zhong and Chen (2021), a well-known information extraction baseline. We finetune PURE over our train set using the default parameters, except for max_span_length, which we set to 40 to accommodate for the longer entities in our data.
Appendix C E2E vs Specialized extraction

### v4

@@ 4.2 Extraction Results
Fine-tuning Mistral-7B on our data yields the best performance across all subtasks. We observe that entity and relation extraction are more challenging than classification for both humans and SOTA LLMs. However, humans still significantly outperform automatic extraction approaches. Appendix B.5 presents an analysis of extraction errors. Interestingly, focusing on a smaller portion of the recombination extraction task is not necessarily easier than performing it end-to-end, as seen in the lower performance of abstract classifiers. We discuss this point further in Appendix B.3.
Large-scale evaluation
To assess extraction quality at scale, we evaluate 2,0002,0002 , 000 CHIMERA examples using a strong LLM-based judge (GPT-4.1). An example is labeled correct if (1) the extracted entities reflect meaningful scientific concepts, and (2) their relation captures a central recombination explicitly described in the abstract. We first validate the judge’s reliability by showing high agreement with human annotations on a representative subset. Applied to the full sample, the judge estimates an extraction accuracy of 80.5580.5580.55%, supporting the robustness of our approach. Notably, most extraction errors are minor, typically involving correct recombinations where the extracted entities are less informative than those in the original abstract (see examples and additional details in Appendix B.7).
Figure 5: Recombination prediction. Given a context string and a query about recombining a graph node, a model trained on CHIMERA suggests plausible recombination directions, leveraging patterns learned from prior examples.
4.3 KB Meta-Science Analysis
@@
Figures 3(a) and 3(b) present the predominant domain pairs for inspiration and blend relations in CHIMERA (above the 0.90.90.9 quantile). The analysis reveals an interesting pattern of a distinct difference in behavior between inspirations and blends: inspirations span a broader range of domains, while blends tend to link within the same or similar domains. This suggests that when human researchers take inspiration they tend to look across more areas other than their own, but tend to look within their own domain when they build approaches by integrating together mechanisms. Inspirations also link more often to areas not covered by the arXiv taxonomy, e.g., cognitive science and zoology. More research building on our initial analysis and KB can shed additional light on the different ways in which scientists combine concepts to form ideas. Table 16 in Appendix D.1 provides a tabular view of this analysis for clarity.
@@
We next analyze how different fields draw inspiration from each other. Figure 3(c) shows the top 101010% cross-domain inspiration sources for three prevalent domains in the graph: cs.RO (Robotics), cs.CV (Computer Vision) and cs.CL (NLP). We observe that while some sources of inspiration (like cognitive-science) are commonly shared across related fields, domains may draw inspiration from unique sources (e.g., from zoology to cs.RO as seen in Figure 1). Interestingly, cs.CV takes more inspiration from cs.CL than vice versa. cs.CL also takes considerably more inspiration from cognitive science than cs.CV, and also takes inspiration from psychology (see example in Table 1), while cs.CV takes more inspiration from biomedical sources. cs.CV also takes inspiration from mathematical topics (discrete math, optimization and control). Appendix B.6 presents examples of such interdisciplinary inspirations.
@@
Table 5: We divide prediction data by the publication years associated with each query (training and validation sets $<$ 2024, test set $\geq$ 2024) to avoid contamination.
Figure 4 shows the percentage of target nodes in domains drawing inspiration from cs.CL (NLP) over five years. We observe two trends: a decrease in intra-domain inspiration (where cs.CL concepts inspire other cs.CL concepts), and an increase in cs.CV (Computer Vision) concepts drawing inspiration from cs.CL.
@@
Table 6: Recombination prediction results. MedR = Median Rank. Fine-tuning on CHIMERA improves MedR 10×. Interestingly, reranking the top-20 answers using RankGPT boosts the H@10 but slightly reduces H@3,5 and MRR.
5 Recombination Prediction
We demonstrate how CHIMERA could be used to train supervised models that recombine concepts and generate novel scientific ideas.
Figure 5 illustrates the recombination prediction task. Given a context string (e.g., “Recent advancements in video generation have struggled to model complex narratives…”) and a query about recombining a graph node (e.g., “What would be a good source of inspiration for video generation?”) the goal is to predict a suitable entity to complete the recombination (e.g., “The concept of storyboarding…”). Formally, given a query with a context string (e.g., a problem, experimental settings, goals), an entity $e$ and a recombination type $\tau$, the task is to predict a different entity $e^{\prime}$ such that $(e,\tau,e^{\prime})$ is a valid edge in CHIMERA.
@@ 5 Recombination Prediction
We start by converting edges to pairs of queries and answers. The queries describe the task inputs: a single graph node, the edge recombination type, and a context string, which we extract from the corresponding abstract using GPT-4o-mini. Note that this process might leak information regarding the answer (the other graph node) into the query. Therefore, we follow it by applying GPT-4o-mini to identify leakages (see examples and implementation details for this step in Appendix E.1). We discard approximately 222222% of pairs due to leaks and split the remainder by publication year, with all papers published after 2024 in the test set. Table 5 summarizes the data splits.
@@
We experiment with zero-shot and finetuned retrievers based on encoders trained before the test set cutoff year (2024). We next explore applying a GPT-4o-based reranker Sun et al. (2023) to the top 20 retrieved results to improve our predictions further. The GPT-4o data cutoff is October 2023, meaning the reranker is also unfamiliar with our test set. Appendix E.2 provides additional implementation details for the prediction baselines.
@@ 5.1 Prediction Results
[…]
Figure 6: Researchers find our recombination suggestions almost as helpful as the gold answer in inspiring ideas, validating our automated evaluation.
@@
We recruit five volunteers with verified research experience (at least one published paper) and assign them examples based on their expertise. Each example includes an inspiration query and suggestions from six sources: (1) Ours: our method, including reranking (2) Gold: the gold answer, (3) Random: a random test-set node, (4) GPT-4o: a GPT-4o generated suggestion, (5) ZS-CHIMERA: zero-shot prediction using our test nodes as candidates, and (6) ZS-SciERC: zero-shot prediction using SciERC-extracted candidates Luan et al. (2018). For baselines returning a ranked list of suggestions, we only use the top result.
Annotators ranked baseline suggestions by their helpfulness in inspiring interesting ideas. Figure 6 reports the median and average rank across 100100100 examples, where lower values indicate better performance. Our approach receives a similar rank as the gold answer, and annotators prefer it to all other baselines. This gives a complementary signal to the automatic evaluation, showing that our recombination prediction approach learns to create helpful recombinations. Appendix F includes further study details and examples of model predictions that participants found especially inspiring.
@@ Conclusions
We present CHIMERA, a novel knowledge base of 28K+ scientist-authored recombinations, capturing how scientists blend concepts and draw inspiration from different areas. CHIMERA supports a wide range of applications—we show its utility for meta-scientific analysis and for fine-tuning models that predict novel, inspiring recombination directions.
@@
As with any automatically-extracted knowledge base, CHIMERA naturally contains some extraction errors (see Appendix B.5). Importantly, our work is the first to explore the new task of extracting recombinations from papers, revealing a gap between the performance of extraction models and humans on the task. As is the case with newly-introduced NLP tasks, future methods trained on our annotated corpus are expected to further improve extraction results, and hence the quality of CHIMERA. However, our analysis shows we already reach good extraction quality overall with minor errors (see Section 4.2), and our downstream applications further demonstrate that the data in CHIMERA can be used to derive utility in scientific meta-analysis and ideation.
Abstract-level scope
CHIMERA focuses on extracting recombination instances from scientific abstracts rather than full papers. This design choice, common in scientific IE tasks Gonzalez et al. (2023); Zhang et al. (2024); Naik et al. (2024), enables more scalable annotation and leverages the fact that abstracts typically summarize key contributions—including conceptual recombinations. In our setting, we focus on capturing cases where a recombination is at the core of a paper’s contribution, hence likely to appear in the abstract. Confirming this intuition, we further conduct an analysis that finds that abstracts cover the vast majority of these cases. Extending extraction methods to full papers could reveal additional recombination patterns in future work.
@@
As in other open-ended creative tasks Jentzsch and Kersting (2023); Meng et al. (2023); Huot et al. (2024), the recombination prediction task admits no single correct answer. Given a problem description, there are many valid ways to blend ideas or draw inspiration, which can lead to false negatives and an overly conservative estimate of model performance. To mitigate this, we conduct a complementary human evaluation. However, due to the expertise required from evaluators, the scale and depth of this assessment are necessarily limited.
@@
Our work leverages a diverse set of models for extraction and prediction, including open-source LLMs (e.g., Mistral-7B, all-mpnet-base), proprietary models (e.g., GPT-4o), and non-generative baselines (e.g., PURE). GPT-4o is used for auxiliary tasks, such as evaluation (judging entity span similarity), analysis (identifying entity’s scientific domain), and to enrich our data (generating a context string for the extracted recombinations). As our primary focus is on building and analyzing the recombination knowledge base, we limit our experiments to these models. Exploring a broader range of models for these auxiliary tasks is an important direction for future work.
Ethical Considerations
To collect human-annotated recombination examples, we recruited crowdworkers through the Upwork platform. All annotators were informed in advance about the nature, purpose, and scope of the annotation task. They were compensated fairly for their time, at rates ranging from \$26 to \$30 per hour. Annotation quality was monitored through overlapping assignments and expert review to ensure reliability and accuracy.
For our human evaluation study, three volunteers with prior research experience participated in ranking model outputs. Participation was entirely voluntary, and no personal or identifying information about the annotators or participants is collected or disclosed.
@@
To support transparency and reproducibility, we release our code, model checkpoints, and the annotated data under an open license. We used AI-based coding assistants (e.g., GitHub Copilot) and language tools for minor code and grammar refinements during development.
Appendix A Annotator Agreement
Following standard practice in information extraction Naik et al. (2023); Sharif et al. (2024), we assess inter-annotator agreement using precision, recall, and F1 scores. Agreement is computed by treating one annotator’s labels as the reference and the other’s as predictions. In addition to measuring entity-level and relation-level agreement, we also evaluate agreement on recombination presence—that is, whether a text expresses a recombination instance, regardless of its type.
We apply the soft entity and relation matching procedure described in Section 4.1 to compute entity and relation agreement. All agreement scores are based on the 49 documents annotated by both annotators (approximately 10% of the full dataset). We treat these agreement measures as a proxy for human-level performance on this task.
A.1 Disagreement Analysis
Annotators’ Disagreement Examples
Abstract: "…This research proposed a framework based on Long Short-Term Memory (LSTM) deep learning network to generate day-ahead hourly temperature forecast… A case study is shown which uses historical in-situ observations and Internet of Things (IoT) observations for New York City, USA. By leveraging the historical air temperature data from in-situ observations, the LSTM model can be exposed to more historical patterns that might not be present in the IoT observations. Meanwhile, by using IoT observations, the spatial resolution of air temperature predictions is significantly improved…"
[…]
Annotator 2: []
Resolution: Upon expert review, Annotator 1’s interpretation was selected, as the authors explicitly describe how the two sources of data serve complementary roles in their method.
Abstract: "…we propose an integrated system that can perform large-scale autonomous flights and real-time semantic mapping in challenging under-canopy environments. We detect and model tree trunks and ground planes from LiDAR data, which are associated across scans and used to constrain robot poses as well as tree trunk models. The autonomous navigation module utilizes a multi-level planning and mapping framework and computes dynamically feasible trajectories that lead the UAV to build a semantic map of the user-defined region of interest in a computationally and storage efficient manner. A drift-compensation mechanism is designed to minimize the odometry drift using semantic SLAM outputs in real time, while maintaining planner optimality and controller stability…"
[…]
Annotator 2: []
Resolution: Annotator 2’s judgment was selected after expert review, as the relation between the two components is not clearly described as a recombination.
Table 7: Examples of annotation disagreements and resolutions by expert review.
To better understand the sources of annotation disagreement, we conducted a qualitative analysis. The most common cause stems from differences in identifying whether a recombination is present at all (see examples in Table 7). In such cases, disagreements were resolved via discussion and expert adjudication. The primary criterion for resolution was whether the authors explicitly describe a recombination as contributing to their approach.
Interestingly, once annotators agreed that a recombination was present, they rarely disagreed on its type, and only a single example exhibited this form of conflict. However, disagreements over which entities the recombination includes were more frequent. These typically fell into two categories:
  1. 1.
    Boundary disagreements, where annotators selected different spans with overlapping meaning. Here, the expert favored the span that preserved more context (e.g., "reinforcement learning which uses traditional time series stock price data" was preferred over "traditional time series stock price data").
  2. 2.
    Conceptual disagreements, where annotators identified fundamentally different entities. These were resolved through further discussion and clarification.
Appendix B Additional Extraction Details
B.1 Recombination keywords
@@
Recombination keywords
combines analogies aggregate intermingle unify blending
combined equivalence aggregation intermingling unification blends
combine equivalent align join weave blend
combination reduction alignment joining weaving blends
combinations reframing amalgamate juxtapose hybrid merge
combining reframe amalgamation juxtaposition merge merges
mixing reformulating assemble link merges unites
mixture casting assembling linkage merging analogy
mix cast associate meld merged analogize
mixed casts association melding conflation analogies
integrates viewing bond mesh couple equivalence
integrating viewed bonding meshing unite equivalent
integrate view bridge perceive unites correlate
integrated inspire bridging perception interplay correlation
connection inspired coalesce relate interconnect envision
synergy inspiration coalescence relation harmonize envisioning
fusion inspires compose splice harmony harmonize
fuses inspiring composition splicing incorporate harmony
unify interconnect incorporation synthesis reduction synthesis
aggregate align inspiring inspire couple conjunction
aggregation reframing inspiration fuse unite conjoin
alignment reframe inspires synthesis
Table 8: Recombination keywords. We use a predefined list of keywords to identify works that are more likely to discuss idea recombination.
B.2 Extraction baselines implementation
@@
We use Mistral-7B as the backbone for our recombination extraction baseline. We fine-tune the model using mistral-finetune [^6] https://github.com/mistralai/mistral-finetune on a single NVIDIA RTX A6000 48GB GPU over 500500500 steps. The training was conducted using the default learning rate of $6.e-5$ and weight decay of 0.10.10.1. We use a batch size of 111 and a maximum sequence length of 409640964096 tokens. mistral-finetune implements Low-Rank Adaptation of LLM (LoRA), a parameter efficient fine-tuning method Hu et al. (2021), which we use with the default rank of 646464. The evaluation uses the corresponding repository, mistral-inference [^7] https://github.com/mistralai/mistral-inference. We rerun the same experiment using Llama-3.1-8B as a backbone, using an additional 500 warm-up steps, a learning rate of $2e-5$ and a weight decay of 0.010.010.01. Figure 7 presents the prompt for these experiments.
@@
In addition to fine-tuning LLMs on our data, we experiment with GoLLIE Sainz et al. (2023), a general IE model fine-tuned to follow any annotation guidelines in a zero-shot fashion. We apply GollIE-13B on our data, using a single NVIDIA RTX A6000 48GB GPU, 1-beam search, and limit the new token number to 128. GoLLIE is finetuned from CODE-LLaMA2, and receives guidelines in the form of data classes describing what objects and properties the model should extract. Figure 8 depicts the guidelines we used to test GoLLIE as an E2E recombination extraction model. In the rare cases where the model returns more than a single recombination type (<10<10< 10), we select the first.
@@
We also experiment with GPT-4o in few-shot settings. We select 45 examples for each example type (blend, inspiration, not-present) from the training data (a total of 135). As Table 2 describes, the training set only has 45 inspiration examples (as opposed to >100>100> 100 blend and not-present examples). 454545 is, therefore, the maximal number of examples per class we can sample while keeping the ICL set balanced. We run each experiment 555 times, sampling a new set of few-shot examples in each, and report the average. Figure 9 presents the prompt for this experiment.
@@
To evaluate entity extraction, we compare our model against GPT-4o in few-shot settings and include 454545 cases per example type, similarly to the E2E experiment. To account for variability due to example selection, we run each experiment 555 times, sampling a new set of few-shot examples in each, and report the average. The total cost of this process sums up to 50\$. The prompt template for this experiment is available on Figure 10.
@@
We experiment with non-generative approaches as well, and compare our model to a SciBERT Zhong and Chen (2021) based token classifier. The encoder uses a standard Hugging-Face implementation of SciBERT, which we train on a single NVIDIA RTX A6000 48GB GPU over 500 steps. We use a weight decay of 0.10.10.1, a learning rate of $6.e-5$ and a batch size of 111. We also experiment with PURE Zhong and Chen (2021), a well-known information extraction baseline. We finetune PURE over our train set using the default parameters, except for max_span_length, which we set to 40 to accommodate for the longer entities in our data.
B.3 E2E vs Specialized extraction